# FoundPose: Unseen Object Pose Estimation with Foundation Features

### v1

@@ 11 Ground-truth instead of CNOS masks
Speed. The presented variants of FoundPose are the most accurate and fastest entries in all three categories evaluated in Tab. 1.111OSOP [82] has a light version that runs at ¡ 1 s per image, but has lower accuracy scores (-3 AR) than the entry in Tab. 1, which takes 5–12 s per image. Notably, the coarse pose estimation stage from FoundPose is significantly faster (1.7 vs 15.5 s) and more accurate than the coarse render-and-compare stage from MegaPose (rows 1 vs 3). The FoundPose variants provide a spectrum of trade-offs between speed and accuracy, with the faster variants (rows 1 and 6) being relevant for online applications such as robotic manipulation, while the more accurate variants being relevant for offline applications such as ground-truth object pose annotation for training supervised methods [89]. The trade-off can be further controlled by, e.g., the template size, the number of PCA components for dimensionality reduction of the patch descriptors, or by different types of the DINOv2 backbone – row 3–4 in Tab. 2 shows that FoundPose runs at around 1.3 s per image if the ViT-L backbone (used for results in Tab. 1) is replaced with the smaller ViT-S. Note that the reported times include also the CNOS segmentation stage that takes around 0.3 s per image.
4.3 Ablation experiments.
Feature extractors. Tab. 2 (rows 1–7) evaluates the performance of FoundPose (without the featuremetric refinement) with different patch descriptors. We achieve the best pose accuracy when we use the output tokens of layer 18 (out of 23) from the DINOv2 ViT-L model with registers[14] as the patch descriptors. We observed also qualitatively that this layer provides a good balance between the positional and semantic information – the resulting correspondences are more geometrically consistent while more discriminative than correspondences obtained using later and earlier layers respectively (Fig. 3).
@@
Furthermore, we evaluate FoundPose with patch descriptors extracted from layer 23 of the SAM ViT-L model [40], from the last backbone layers of the feature matching pipelines LoFTR [86] and S2DNet [21], from the last layer of the CLIP image encoder [75], and with patch descriptors defined by the classical SIFT descriptor [52] calculated on a regular 2D grid with the cell size of 7 px. In case of S2DNet, patch descriptors were obtained by sampling the last CNN feature map. All method parameters were fixed in these experiments, only the patch descriptors were different. As shown in Tab. 2, patch descriptors from DINOv2 are the key enabler of FoundPose, yielding significantly higher accuracy than the alternatives (we only show performance of the better alternatives).
@@
Template retrieval. Next, on rows 8–9 in Tab. 2, we evaluate the proposed bag-of-words template retrieval approach against an alternative based on matching the cls token from DINOv2. In this alternative, we use the cls token from layer 18 of DINOv2 ViT-L as the template & crop descriptor, and retrieve templates whose descriptors have the highest cosine similarity with the descriptor of the query image crop. A similar approach is used in ZS6D [3] for initial viewpoint estimation and in CNOS [60] for object identification. Compared to the bag-of-words descriptor constructed from patch descriptors, we observed that the cls token is less robust to occlusions and contains a limited information about the object pose (as also shown in Fig. 5 in [60]). To avoid influence of the background on the cls token, we additionally evaluate a variant where pixels outside the object mask are made black, as in [3, 60]. As shown on rows 8 vs 9 in Tab. 2, this modification improves the performance of the cls-based approach, but is still far from the performance of our bag-of-words approach (row 1). When tried with the cls token from layer 23, the alternative retrieval approach achieved even lower accuracy.
@@
We further analyze the quality of templates retrieved by the bag-of-words approach by evaluating the object pose associated with the top retrieved template. The pose from the template is adjusted such as the 2D bounding circle of the object is aligned with the 2D bounding circle of the object segmentation mask in the query image. This approach is surprisingly effective, reaching accuracy close to the coarse poses from MegaPose while being 15 times faster (row 10 in Tab. 2 vs row 3 in Tab. 1).
@@
Ground-truth segmentation masks. A common source of FoundPose failure cases are erroneous object segmentation masks. When the ground-truth masks are used instead of the masks predicted by CNOS [60], FoundPose achieves a large improvement of +6–19 AR (rows 1 vs 11 in Tab. 2).
@@ 5 Conclusion
We have proposed a practical method for model-based pose estimation of unseen objects which demonstrates the effectiveness of features from the DINOv2 foundation model for this task. The method can handle a broad range of objects, including texture-less and symmetric objects, can scale to large numbers of objects thanks to a minimal object representation, and is computationally efficient, mostly thanks to the bag-of-words template retrieval which simplifies the subsequent process of establishing 2D-3D correspondences by providing a small set of reliable templates. Without any object-specific training, variants of FoundPose outperform existing RGB methods on the BOP benchmark.

### v2

@@ 4.2 Main results
Speed. The presented variants of FoundPose are the most accurate and the second-fastest entries in all three categories evaluated in Tab. 1.111The time of GigaPose [59], the only faster method, was obtained with a more powerful GPU (V100 48GB vs P100 16GB used in our experiments). [^2] OSOP [75] has a light version that runs at ¡ 1 s per image, but has lower accuracy scores (-3 AR) than the entry in Tab. 1, which takes 5–12 s per image. Notably, the coarse pose estimation stage from FoundPose is significantly faster (1.7 vs. 15.5 s) and more accurate than the coarse render-and-compare stage from MegaPose (rows 1 vs. 4). The presented FoundPose variants provide a spectrum of trade-offs between speed and accuracy, with the faster variants (rows 1 and 7) being relevant for online applications such as robotic manipulation, while the more accurate variants being relevant for offline applications such as ground-truth object pose annotation for training supervised methods [82]. The trade-off can be further controlled by, e.g., the template size or different types of the DINOv2 backbone – rows 3–4 in Tab. 2 show that FoundPose runs at around 1.3 s per image if the ViT-L backbone (used for results in Tab. 1) is replaced with the smaller ViT-S. Note that the reported times also include the CNOS segmentation stage that takes around 0.3 s per image. The offline object onboarding takes less than 5 minutes, as requested by BOP [34].
4.3 Ablation experiments
@@
Feature extractors. Tab. 2 (rows 1–7) evaluates the performance of FoundPose (without the featuremetric refinement) with different patch descriptors. We achieve the best pose accuracy when we use the output tokens of layer 18 (out of 23) from the DINOv2 ViT-L model with registers[15] as the patch descriptors. We also observed qualitatively that this layer provides a good balance between the positional and semantic information – the resulting correspondences are more geometrically consistent than correspondences obtained using later layers (Fig. 3). Furthermore, we evaluate FoundPose with patch descriptors extracted from layer 23 of SAM ViT-L [40], from the last backbone layers of the feature matching pipelines LoFTR [78] and S2DNet [23], from the last layer of the CLIP image encoder [70], and with patch descriptors defined by the classical SIFT descriptor [50] calculated on a regular 2D grid with the cell size of 7 px. In the case of S2DNet, descriptors were obtained by sampling the last CNN feature map. All method parameters were fixed in these experiments, only the patch descriptors differed. As shown in Tab. 2, patch descriptors from DINOv2 are the key enabler of FoundPose, yielding significantly higher accuracy than the alternatives (we only show the performance of the better alternatives).
@@
Template retrieval. Next, on rows 8–9 in Tab. 2, we evaluate the proposed bag-of-words template retrieval approach against an alternative based on matching the cls token from DINOv2, which is used in CNOS [58] for object identification. In this alternative, we use the cls token from layer 18 of DINOv2 ViT-L as the template and crop descriptor and retrieve templates whose descriptors have the highest cosine similarity with the crop descriptor. Compared to the bag-of-words descriptor constructed from patch descriptors, we observed that the cls token is less robust to occlusions and contains limited information about the object pose (as also shown in Fig. 5 of [58]). To avoid the influence of the background on the cls token, we additionally evaluate a variant where pixels outside the object mask are made black, as in [3, 58]. As shown on rows 8 vs. 9 in Tab. 2, this modification improves the performance of the cls-based approach but is still far from the performance of our bag-of-words approach (row 1). When tried with the cls token from layer 23, the alternative retrieval approach achieved even lower accuracy.
@@
We further analyze the quality of templates retrieved by the bag-of-words approach by evaluating the object pose associated with the top retrieved template. The pose from the template is adjusted such that the 2D bounding circle of the object is aligned with the 2D bounding circle of the object segmentation mask in the query image. This approach is surprisingly effective, reaching accuracy close to the coarse poses from MegaPose while being 15 times faster (row 10 in Tab. 2 vs row 4 in Tab. 1).
Another possible retrieval approach is to directly match patches of each query-template pair, and calculate the pair similarity from the matching quality, e.g., as done by Goodwin et al. [24]. This approach needs a kNN search for each pair and is noticeably less efficient than our approach (0.64s vs 0.0008s per query; measured with faiss[38]).
@@
Ground-truth segmentation masks. A common source of failure cases are erroneous segmentation masks. When ground-truth masks are used instead of masks predicted by CNOS [58], FoundPose achieves a large +6–19 AR improvement (rows 1 vs 11 in Tab. 2).
Effect of other parameters. We also evaluated FoundPose with different numbers of templates (400, 800, 1600), PCA components (128, 256, 512, 1024), visual words (1024, 2048, 4096), and RANSAC iterations (200, 400, 800). These settings influence memory requirements but lead to only marginal differences in speed and accuracy.
[…]
@@ 5 Conclusion
We have proposed an RGB method for model-based pose estimation of unseen objects, which significantly outperforms existing methods on the standard BOP benchmark. We believe that achieving this without any object- nor task-specific training, just with a frozen vision foundation model, is an important and non-obvious outcome. Furthermore, we have shown that the method can be seamlessly combined with an existing render-and-compare refinement approach to achieve RGB-only state-of-the-art results. Our strong results are encouraging to revisit efficient classical computer vision which is often overlooked in the modern literature.